OpenStudy (anonymous):

can someone help me begin the problem 4 x log2(x-1)=3?

OpenStudy (anonymous):



OpenStudy (mertsj):

Divide both sides by 4

OpenStudy (mertsj):

\[\log_{2}(x-1)=\frac{3}{4} \]

OpenStudy (mertsj):

\[2^{\frac{3}{4}}=x-1\]

OpenStudy (mertsj):

\[x=1+2^{\frac{3}{4}}=2.68\]

OpenStudy (anonymous):

did you move the 2 from the log and put it by the 3/4?

OpenStudy (anonymous):

never mind, I see what you did. Thank you!

OpenStudy (mertsj):

yw